# Generalizing subdiffusive Black-Scholes model by variable exponent: Model transformation and numerical approximation

### v1

@@
Abstract
This work generalizes the subdiffusive Black-Scholes model by introducing the variable exponent in order to provide adequate descriptions for the option pricing, where the variable exponent may account for the variation of the memory property. In addition to standard nonlinear-to-linear transformation, we apply a further spatial-temporal transformation to convert the model to a more tractable form in order to circumvent the difficulties caused by the “non-positive, non-monotonic” variable-exponent memory kernel. An interesting phenomenon is that the spatial transformation not only eliminates the advection term but naturally turns the original noncoercive spatial operator into a coercive one due to the specific structure of the Black-Scholes model, which thus avoids imposing constraints on coefficients. Then we perform numerical analysis for both the semi-discrete and fully discrete schemes to support numerical simulation. Numerical experiments are carried out to substantiate the theoretical results.
keywords:
Black-Scholes model , Subdiffusion , Variable exponent , Error estimate , Option pricing
journal: Economic Modelling
@@ 1 Introduction
$^{R}\partial^{\alpha}_{t}g(t):=\partial_{t}\int_{t}^{T}\frac{g(s)-g(T)}{\Gamma(1-\alpha)(s-t)^{\alpha}}ds,$ (1.2)
and some progresses on the resulting subdiffusive Black-Scholes model have been reached in the past few decades [6, 7, 8, 9, 11, 12, 16, 17, 18, 19, 22, 23]. Nevertheless, a constant exponent may not suffice to accommodate the variation of the memory property caused by, e.g. the uncertainties or fluctuations in the financial market. A possible remedy is to introduce the variable exponent. In [13], a variable-exponent fractional option pricing model is applied to accommodate various phenomena in financial activities such as the seasonal changes and temporary crises, and the effectiveness of the variable-exponent model has been demonstrated based on the real option price data. Thus, we consider the following subdiffusive Black-Scholes model with variable exponent $10<\alpha(t)<$ 1 and the expiry $T>0$
@@
Here $c_{l}$, $c_{r}$, $c_{t}$ are given data and ${}^{R}\partial^{\alpha(t)}_{t}$ with $10<\alpha(t)<$ 1 is the variable-exponent fractional differential operator [14]
$^{R}\partial^{\alpha(t)}_{t}g(t):=\partial_{t}\int_{t}^{T}\frac{g(s)-g(T)}{\Gamma(1-\alpha(s-t))(s-t)^{\alpha(s-t)}}ds.$ (1.4)
As the exact solutions of model (1.3) are in general not available, it is necessary to study numerical methods. Extensive numerical results for model (1.3) (or its variant (2.6) by standard logarithmic transformation) with $\alpha(t)\equiv\alpha$ for some constant $10<\alpha<$ 1 have been obtained, while there are rare studies for the variable-exponent case. A main difficulty is that the variable-exponent memory kernel is non-positive and non-monotonic such that conventional numerical analysis methods do not apply. In a recent work [24], a relevant variable-exponent fractional Black-Scholes model is considered, where an additional term $\partial_{t}v$ is included in (1.3). Due to this additional leading term, the impact of the variable exponent is significantly weakened such that the numerical analysis could be performed.
@@ 2.1 Nonlinear-to-linear transformation
[…]
Therefore, if we employ the transformation $x=\ln S$ and accordingly define $u(x,t)=v(e^{x},T-t)$, the subdiffusive Black-Scholes model (1.3) could be transformed into the following form
@@
where $\bar{a}:=\ln a$, $\bar{b}:=\ln b$, $\bar{c}_{l}(t):=c_{l}(T-t)$, $\bar{c}_{r}(t):=c_{r}(T-t)$, $\bar{c}_{t}(x):=c_{t}(e^{x})$ and $\partial^{\alpha(t)}_{t}$ is defined by
[…]
@@ 2.2 Further spatial-temporal transformation
According to equation (2.5) and the transformed technique in [25], we perform the following convolution with equation (2.7)
$\beta_{\alpha_{0}}*\left[(k*\partial_{t}\phi)-\frac{1}{2}\sigma^{2}\partial_{x}^{2}\phi+\lambda\phi-\chi\right]=0,$
@@ 2.3 Assumptions
For a positive integer $m$, a real number $1\leq p\leq\infty$ and an interval $\mathcal{I}$, let $W^{m,p}({\mathcal{I}})$ be the Sobolev space of functions with weak derivatives up to order $m$ in $L^{p}(\mathcal{I})$, where $L^{p}(\mathcal{I})$ refers to the space of $p$th Lebesgue integrable functions on $\mathcal{I}$. Let $H^{m}({\mathcal{I}}):=W^{m,2}({\mathcal{I}})$ and $H^{m}_{0}({\mathcal{I}})$ be its subspace with the zero boundary conditions up to order $m-1$. For a Banach space ${\cal X}$, let $W^{m,p}(\mathcal{I};\mathcal{X})$ be the space of functions in $W^{m,p}(\mathcal{I})$ with respect to $\|\cdot\|_{\cal X}$. All spaces are equipped with the standard norms [1]. We set $\|\cdot\|:=\|\cdot\|_{L^{2}({\mathcal{I}})}$ for brevity if $\mathcal{I}$ represents a spatial interval, and drop the notation ${\mathcal{I}}$ in the spaces and norms if $\mathcal{I}$ represents a spatial interval, e.g. we denote $\|\cdot\|_{H^{2}}=\|\cdot\|_{H^{2}(\mathcal{I})}$.
Throughout this work, we consider the smooth variable exponent, i.e., we assume that $10<\alpha(t)\leq\alpha^{*}<$ 1 for $t\in[0,T]$ and $\alpha\in W^{1,\infty}(0,T)$. Furthermore, we use $C$ to denote a generic positive constant that may assume different values at different occurrences.
Following the regularity results of [25, Theorems 4.2-4.3], the solutions to the model (2.15)-(2.16) satisfy the following estimates under sufficiently regular data
@@
Finally, we impose the condition $\alpha^{\prime}(0)=0$ throughout the work. It has been demonstrated in [26, Section 3.1] that this constraint can be arbitrarily weak. Specifically, for any smooth function $\alpha(t)$, one could construct a sequence of smooth functions $\{\alpha_{\sigma}(t)\}_{\sigma>0}$ satisfying $\alpha_{\sigma}^{\prime}(0)=0$ such that $\max_{t\in[0,T]}|\alpha(t)-\alpha_{\sigma}(t)|\rightarrow 0$ as $\sigma\rightarrow 0$.
@@
Then we yield $|\partial_{t}D(t\varsigma)|\leq C$ with $t\in[0,T]$. We finish the proof of the lemma. ∎
@@ 3.1 Establishment of time-discrete scheme
$\omega(t_{n})+(q^{\prime}*\omega)(t_{n})-\frac{1}{2}\sigma^{2}\beta_{\alpha_{0}}*\partial_{x}^{2}\omega(t_{n})+\lambda\beta_{\alpha_{0}}*\omega(t_{n})=F(t_{n}).$ (3.1)
@@ 3.2 Analysis of time-discrete scheme
where the norm $\|\cdot\|_{A}$ defined in (3.18).
Proof.
@@
$\max\limits_{1\leq n\leq N}\|u(t_{n})-U^{n}\|_{A}\leq C(T)\tau^{\frac{1}{2}+\frac{3}{2}\alpha_{0}}.$
Proof.
@@
which finishes the proof of Theorem 3.2. ∎
@@ 4 Fully discrete finite element scheme
Define a quasi-uniform partition of $[\bar{a},\bar{b}]$ with mesh diameter $h$ and let $S_{h}$ be the space of continuous and piecewise linear functions on $[\bar{a},\bar{b}]$ with respect to the partition. Let $I$ be the identity operator. The Ritz projection $\mathrm{\Pi}_{h}:H_{0}^{1}([\bar{a},\bar{b}])\to S_{h}([\bar{a},\bar{b}])$ defined by $(\nabla(\psi-\mathrm{\Pi}_{h}\psi),\nabla\varrho)=0$ for any $\varrho\in S_{h}$ has the approximation property [20]
$\begin{array}[]{lll}\|(I-\mathrm{\Pi}_{h})\psi\|\leq Ch^{2}\|\psi\|_{H^{2}},\quad\forall\psi\in H^{2}([\bar{a},\bar{b}])\cap H_{0}^{1}([\bar{a},\bar{b}]).\end{array}$ (4.1)
Multiply equation (3.12) by $\Lambda\in H_{0}^{1}[\bar{a},\bar{b}])$ on $[\bar{a},\bar{b}]$ to obtain the weak formulation for any $\Lambda\in H_{0}^{1}([\bar{a},\bar{b}])$ and $n=1,2,\dots,N$
@@
$=(F^{n},\Lambda)+(R^{n},\Lambda),\quad 1\leq n\leq N,$ (4.2)
Drop the local truncation error terms in (4.2) to obtain a fully-discrete finite element scheme: find $\mathcal{W}^{n}\in S_{h}$ for $1\leq n\leq N$ such that
$(\mathcal{W}^{n},\Lambda)+(\sum\limits_{j=1}^{n}\varpi_{n,j}\mathcal{W}^{j},\Lambda)$ $+\lambda(\sum\limits_{j=1}^{n}\widetilde{\varpi}_{n,j}\mathcal{W}^{j},\Lambda)+\frac{1}{2}\sigma^{2}(\sum\limits_{j=1}^{n}\widetilde{\varpi}_{n,j}\nabla\mathcal{W}^{j},\nabla\Lambda)$
@@
After solving the numerical solution $\mathcal{W}^{n}$ in (4.3), we further define the numerical solution $U^{n}$ of equation (2.6) as follows
$U^{n}=\exp\left((\frac{1}{2}-\frac{r}{\sigma^{2}})x\right)(\mathcal{W}^{n}+\bar{c}_{t}^{*}).$
@@ 4.1 Error estimate of fully discrete scheme
Assume that $\omega$ satisfies the regularity assumption in (2.18) and $\mathcal{W}^{n}$ be the numerical approximation of $\omega$. Then, the following convergence result holds with $T<\infty$
$\sqrt{\tau\sum\limits_{n=1}^{N}\|\omega^{n}-\mathcal{W}^{n}\|^{2}}\leq C(T)(\tau^{\frac{1}{2}+\frac{3}{2}\alpha_{0}}+h^{2}).$ (4.4)
Proof.
Set $\omega^{n}-\mathcal{W}^{n}=\xi^{n}+\eta^{n}$ with $\eta^{n}:=\omega^{n}-\rm\Pi_{h}\omega^{n}$ and $\xi^{n}:=\rm\Pi_{h}\omega^{n}-\mathcal{W}^{n}$. Then we subtract (4.3) from (4.2) and select $\Lambda=\xi^{n}$ to obtain
@@
We combine this with $\|\omega^{N}-\mathcal{W}^{N}\|\leq Ch^{2}$ to get
$\begin{split}\|\omega^{N}-\mathcal{W}^{N}\|_{B}^{2}\leq C(T)^{2}(\tau^{\frac{1}{2}+\frac{3}{2}\alpha_{0}}+h^{2})^{2}.\end{split}$
The proof is completed by using (4.9). ∎

### v2

# Corresponding author

@@
[…]
Corresponding author
[…]
1. School of Statistics and Mathematics, Shandong University of Finance and Economics, Jinan 250014, China
E-mail :: zmh_\_1212@163.com; liuyaxue@sdufe.edu.cn
2. School of Mathematics and Statistics, Hunan Normal University, Changsha, Hunan 410081, China
E-mail :: liumengmeng423@163.com
3. School of Mathematics, Shandong University, Jinan 250100, China
E-mail :: wlqiu@sdu.edu.cn
4. School of Mathematics, State Key Laboratory of Cryptography and Digital Economy Security, Shandong University, Jinan 250100, China
E-mail :: xzheng@sdu.edu.cn
GENERALIZING SUBDIFFUSIVE BLACK-SCHOLES MODEL BY VARIABLE EXPONENT: MODEL TRANSFORMATION AND NUMERICAL APPROXIMATION
Meihui ZHANG1 Yaxue LIU1 Mengmeng LIU2,3∗ Wenlin QIU3 Xiangcheng ZHENG4
[…]
@@ 1 Introduction
$^{R}\partial^{\alpha}_{t}g(t):=\partial_{t}\int_{t}^{T}\frac{g(s)-g(T)}{\Gamma(1-\alpha)(s-t)^{\alpha}}\mathrm{d}s,$ (1.2)
and some progresses on the resulting subdiffusive Black-Scholes model have been reached in the past few decades [19, 22, 23, 25, 28, 29, 30]. Nevertheless, a constant exponent may not suffice to accommodate the variation of the memory property caused by, e.g. the uncertainties or fluctuations in the financial market [9, 10, 11, 13, 14, 17, 24]. A possible remedy is to introduce the variable exponent. In [18], a variable-exponent fractional option pricing model is applied to accommodate various phenomena in financial activities such as the seasonal changes and temporary crises, and the effectiveness of the variable-exponent model has been demonstrated based on the real option price data. Thus, we consider the following subdiffusive Black-Scholes model with variable exponent $0<\alpha(t)<1$ and the expiry $T>0$
@@
Here $c_{l}$, $c_{r}$, $c_{t}$ are given data and $^{R}\partial^{\alpha(t)}_{t}$ with $0<\alpha(t)<1$ is the variable-exponent fractional differential operator [20]
$^{R}\partial^{\alpha(t)}_{t}g(t):=\partial_{t}\int_{t}^{T}\frac{g(s)-g(T)}{\Gamma(1-\alpha(s-t))(s-t)^{\alpha(s-t)}}\mathrm{d}s.$ (1.4)
As the exact solutions of model (1.3) are in general not available, it is necessary to study numerical methods. Extensive numerical results for model (1.3) (or its variant (2.6) by standard logarithmic transformation) with $\alpha(t)\equiv\alpha$ for some constant $0<\alpha<1$ have been obtained, while there are rare studies for the variable-exponent case. A main difficulty is that the variable-exponent memory kernel is non-positive and non-monotonic such that conventional numerical analysis methods do not apply. In a recent work [31], a relevant variable-exponent fractional Black-Scholes model is considered, where an additional term $\partial_{t}v$ is included in (1.3). Due to this additional leading term, the impact of the variable exponent is significantly weakened such that the numerical analysis could be performed.
@@ 2.1 Nonlinear-to-linear transformation
$\begin{array}[]{rl}{}^{R}\partial^{\alpha(t)}_{t}g(t)&\displaystyle=\partial_{t}\int_{t}^{T}\frac{g(s)-g(T)}{\Gamma(1-\alpha(s-t))(s-t)^{\alpha(s-t)}}\mathrm{d}s\\[10.84006pt] &\displaystyle=-\partial_{\tau}\int_{T-\tau}^{T}\frac{g(s)-g(T)}{\Gamma(1-\alpha(s-T+\tau))(s-(T-\tau))^{\alpha(s-T+\tau)}}\mathrm{d}s\\[10.84006pt] &\displaystyle={-\partial_{\tau}}\int_{0}^{\tau}\frac{g(T-\theta)-g(T)}{\Gamma(1-\alpha(\tau-\theta))(\tau-\theta)^{\alpha(\tau-\theta)}}\mathrm{d}\theta.\end{array}$ (2.1)
Then we intend to employ the transformation $x=\ln S$ and accordingly define $u(x,t)=v(e^{x},T-t)$. For the typical case where $S\in(a,b)=(0,\infty)$, the logarithmic transformation will lead to an unbounded domain. In this case, we follow the standard treatment, see e.g. [13, 30], to truncate the unbounded domain into a finite interval. Thus, the subdiffusive Black-Scholes model (1.3) could be transformed into the following form
@@
where $(\bar{a},\bar{b})$ is defined as $(\ln a,\ln b)$ if $(\ln a,\ln b)$ is bounded or the truncated interval if $(\ln a,\ln b)$ is unbounded, $\bar{c}_{l}(t):=c_{l}(T-t)$, $\bar{c}_{r}(t):=c_{r}(T-t)$, $\bar{c}_{t}(x):=c_{t}(e^{x})$ and $\partial^{\alpha(t)}_{t}$ is defined by
$\partial^{\alpha(t)}_{t}u(x,t):=\partial_{t}\int_{0}^{t}\frac{u(x,s)-u(x,0)}{\Gamma(1-\alpha(t-s))(t-s)^{\alpha(t-s)}}\mathrm{d}s.$ (2.3)
@@ 2.2 Further spatial-temporal transformation
According to equation (2.5) and the convolution method proposed in [32], we perform the following convolution with equation (2.7)
$\beta_{\alpha_{0}}*\left[(k*\partial_{t}\phi)-\frac{1}{2}\sigma^{2}\partial_{x}^{2}\phi+\lambda\phi-\chi\right]=0,$ (2.11)
@@
Remark 2.1.
This work considers the case that the volatility $\sigma$ is a constant. Although the nonlinear-to-linear transformation in Section 2.1 works for the more general case $\sigma=\sigma(S,t)$, the spatial-temporal transformation in this section will no longer work. A possible remedy for the case $\sigma=\sigma(t)$ is to replace the convolution method used in (2.11) by the perturbation method proposed in [32], which only performs a kernel splitting for $k$ without affecting other terms. We will investigate this more general case in the future.
@@ 2.3 Assumptions
For a positive integer $m$, a real number $1\leq p\leq\infty$ and an interval $\mathcal{I}$, let $W^{m,p}({\mathcal{I}})$ be the Sobolev space of functions with weak derivatives up to order $m$ in $L^{p}(\mathcal{I})$, where $L^{p}(\mathcal{I})$ refers to the space of $p$th Lebesgue integrable functions on $\mathcal{I}$. Let $H^{m}({\mathcal{I}}):=W^{m,2}({\mathcal{I}})$ and $H^{m}_{0}({\mathcal{I}})$ be its subspace with the zero boundary conditions up to order m1m-1. For a Banach space $\cal X$}, let $W^{m,p}(\mathcal{I};\mathcal{X})$ be the space of functions in $W^{m,p}(\mathcal{I})$ with respect to $\|\cdot\|_{\cal X}$. All spaces are equipped with the standard norms [1]. We set $\|\cdot\|:=\|\cdot\|_{L^{2}({\mathcal{I}})}$ for brevity if $\mathcal{I}$ represents a spatial interval, and drop the notation $\mathcal{I}$} in the spaces and norms if $\mathcal{I}$ represents a spatial interval, e.g. we denote $\|\cdot\|_{H^{2}}=\|\cdot\|_{H^{2}(\mathcal{I})}$.
Throughout this work, we consider the smooth variable exponent, i.e., we assume that $0<\alpha(t)\leq\alpha^{*}<1$ for $t\in[0,T]$ and $\alpha\in W^{1,\infty}(0,T)$. Furthermore, we use $C$ to denote a generic positive constant that may assume different values at different occurrences.
Following the regularity results of [32, Theorems 4.2-4.3], we assume that the solutions to the model (2.16)-(2.17) satisfy the following estimates under sufficiently regular data
@@
Finally, we impose the condition $\alpha^{\prime}(0)=0$ throughout the work. It has been demonstrated in [33, Section 3.1] that this constraint can be arbitrarily weak. Specifically, for any smooth function $\alpha(t)$, one could construct a sequence of smooth functions $\{\alpha_{\sigma}(t)\}_{\sigma>0}$ satisfying $\alpha_{\sigma}^{\prime}(0)=0$ such that $\max_{t\in[0,T]}|\alpha(t)-\alpha_{\sigma}(t)|\rightarrow 0$ as $\sigma\rightarrow 0$.
@@
Then we yield $|\partial_{t}D(t\varsigma)|\leq C$ with $t\in[0,T]$. We finish the proof of the lemma.
@@ 3.1 Establishment of time-discrete scheme
$\omega(t_{n})+(q^{\prime}*\omega)(t_{n})-\frac{1}{2}\sigma^{2}c(\beta_{\alpha_{0}}*\partial_{x}^{2}\omega)(t_{n})+\lambda\beta_{\alpha_{0}}*\omega(t_{n})=F(t_{n}).$ (3.1)
@@
Since $\mathcal{W}^{n}$ is a solution to (3.13)-(3.14), we can directly obtain a numerical solution to (2.6) through the relation $U^{n}=\exp\left((\frac{1}{2}-\frac{r}{\sigma^{2}})x\right)(\mathcal{W}^{n}+\bar{c}_{t}^{*})$.
@@ 3.2 Analysis of time-discrete scheme
where the norm $\|U^{N}\|_{A}:=\sqrt{\tau\sum\limits_{l=1}^{N}\|U^{l}\|^{2}}$, and $\tau$ sufficiently small.
Proof 3.2.
@@
$\|u(t_{N})-U^{N}\|_{A}\leq C(T)\tau^{\frac{1}{2}+\frac{3}{2}\alpha_{0}}.$
Proof 3.4.
@@
which finishes the proof of Theorem 3.3.
@@ 4 Fully discrete finite element scheme
Define a quasi-uniform partition of $\bar{a},\bar{b}]$ with mesh diameter $h$ and let $S_{h}$ be the space of continuous and piecewise linear functions on $\bar{a},\bar{b}]$ with respect to the partition. Let $I$ be the identity operator. The Ritz projection $\mathrm{\Pi}_{h}:H_{0}^{1}(\bar{a},\bar{b})\to S_{h}$ defined by $(\nabla(\psi-\mathrm{\Pi}_{h}\psi),\nabla\varrho)=0$ for any $\varrho\in S_{h}$ has the approximation property [26]
$\begin{array}[]{lll}\|(I-\mathrm{\Pi}_{h})\psi\|\leq Ch^{2}\|\psi\|_{H^{2}},\quad\forall\psi\in H^{2}(\bar{a},\bar{b})\cap H_{0}^{1}(\bar{a},\bar{b}).\end{array}$ (4.1)
Multiply equation (3.12) by $\Lambda\in H_{0}^{1}[\bar{a},\bar{b}])$ on $\bar{a},\bar{b}]$ to obtain the weak formulation for any $\Lambda\in H_{0}^{1}([\bar{a},\bar{b}])$ and $n=1,2,\dots,N$
@@
$=(F^{n},\Lambda)+(R^{n},\Lambda),\quad 1\leq n\leq N.$ (4.2)
Drop the local truncation error terms in (4.2) to obtain a fully-discrete finite element scheme: find $\mathcal{W}_{h}^{n}\in S_{h}$ for $1\leq n\leq N$ such that
$(\mathcal{W}_{h}^{n},\Lambda)+(\sum\limits_{j=1}^{n}\varpi_{n,j}\mathcal{W}^{j},\Lambda)$ $+\lambda(\sum\limits_{j=1}^{n}\widetilde{\varpi}_{n,j}\mathcal{W}_{h}^{j},\Lambda)+\frac{1}{2}\sigma^{2}(\sum\limits_{j=1}^{n}\widetilde{\varpi}_{n,j}\nabla\mathcal{W}_{h}^{j},\nabla\Lambda)$
@@
After solving the numerical solution $\mathcal{W}_{h}^{n}$ in (4.3), we further define the numerical solution $U_{h}^{n}$ of equation (2.6) as follows
$U_{h}^{n}=\exp\left((\frac{1}{2}-\frac{r}{\sigma^{2}})x\right)(\mathcal{W}_{h}^{n}+\bar{c}_{t}^{*}).$
@@ 4.1 Error estimate of fully discrete scheme
We now estimate the error of the full discrete scheme (4.3) in the following theorem.
@@
Assume that $\omega$ satisfies the regularity assumption in (2.19) and $\mathcal{W}_{h}^{n}$ be the numerical approximation of $\omega$. Then, the following convergence result holds with $T<\infty$ and $\tau$ sufficiently small
$\sqrt{\tau\sum\limits_{n=1}^{N}\|\omega^{n}-\mathcal{W}_{h}^{n}}\|^{2}\leq C(T)(\tau^{\frac{1}{2}+\frac{3}{2}\alpha_{0}}+h^{2}).$ (4.4)
Proof 4.2.
Set $\omega^{n}-\mathcal{W}_{h}^{n}=\xi^{n}+\eta^{n}$ with $\eta^{n}:=\omega^{n}-\rm\Pi_{h}\omega^{n}$ and $\xi^{n}:=\rm\Pi_{h}\omega^{n}-\mathcal{W}_{h}^{n}$. Then we subtract (4.3) from (4.2) and select $\Lambda=\xi^{n}$ to obtain
@@
We combine this with $\|\omega^{N}-\mathcal{W}_{h}^{N}\|\leq Ch^{2}$ to get
$\begin{split}\|\omega^{N}-\mathcal{W}_{h}^{N}\|_{B}^{2}\leq C(T)^{2}(\tau^{\frac{1}{2}+\frac{3}{2}\alpha_{0}}+h^{2})^{2}.\end{split}$
The proof is completed by using (4.9).